### v2

@@ Abstract
In this paper I develop and characterize a random attention model with unobserved reference alternatives. The decision maker pays attention to different subsets of the available set of alternatives randomly. The reference alternatives are exactly those alternatives that are always paid attention to, i.e. they are attention-privileged and these alternatives are supposed to be unknown to the outside observer. The characterization allows for a complete identification of the reference alternatives and a coarse identification of the underlying preferences. I then restrict the model by considering the independent random attention function and provide a complete identification of the underlying preferences as well.
@@ 1 Introduction
Decision-making is often constrained by limited attention, whether due to limits to cognitive processing because of too many alternatives to choose from or due to lack of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times under the same primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices with the main research task then been to ascertain the likely underlying heuristic and the primitives used by the DM. Cattaneo et al. (2020) introduced the framework of random attention models (RAM) to deal with the analysis of stochastic choice functions with random choice due to limited attention.
@@
There has been plenty of literature dealing with observable status quo affecting the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. These authors were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. It is natural to presume that a reference alternative can direct attention when the underlying heuristic is applied due to limited attention, for which they provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
@@
In certain cases, it may make sense that the reference is observable. For instance, for a decision maker in the market for a new car, his current car can be a reference. However, in many situations, it is not quite realistic to assume the knowledge of the DM’s reference to anyone else but the DM himself. Secondly, to analyse a choice problem with observable references, we have to deal with a family of stochastic choice functions, one for each alternative as the reference, which is quite a large data requirement. If we restrict the references to be unknown we deal with only menus and stochastic choice functions over alternatives and menus. We do not have to deal with families of stochastic choice functions for each possible reference.
@@
It is also not obvious that a reference alternative has to be unique. In what follows we allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged. However, since I am looking at choices from menus that depend on reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below two examples that match the mentioned assumptions:
  • Example 1: Michael Jordan is associated with the Air Jordan brand, a fan of him who is interested in buying shoes might consider an Air Jordan shoe as a reference.
@@
  • Example 2: Suppose your neighbor buys a new flat-screen 4K TV and invites you over to watch a football game together. You are startled by the picture quality and ask about the product. The prior knowledge of this particular TV will mean that this product will act as a reference point the next time you buy a TV.
In the above examples, it is not obvious that the reference will be observable to the outside observer. Neither is there any evidence that the reference is unique, a person could be a fan of multiple athletes each with their own clothing line, or have multiple suggestions when deciding to buy a new TV.
@@
In this paper, I look at a random attention model with unobservable menu-independent reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the RAM-UR by utilizing three axioms and identifying the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences) We can uniquely identify the set of references but the underlying preferences are only coarsely identified, i.e. the identification is with respect to the relative ranking between any alternative and any reference alternative and the transitive closure thereof. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be fully identified. The existence of menus with no references is crucial in making the preference identification in this case which is possible here because of the menu-independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive which precludes them from considering menus without references. I also elaborate on the fact that rational choice is a boundary case of the RAM-UR as well as the RAM-UR-IRA, and the IRA model is also a boundary case of the RAM-UR-IRA [^3] Boundary cases can be of two types: When the set of references is the set of all alternatives, or when the set of references is empty..
@@
The rest of the paper is organized as follows. In Section 2, I introduce the general RAM-UR model, axiomatically characterize it, and discuss the identification results. The characterization of the RAM-UR-IRA is then provided in Section 3, where I also discuss the nestedness properties of the corresponding random attention models. Section 4 relates my work to the existing literature. All proofs are relegated to Appendix 1.
@@ 2.1 Primitives
Below I abuse the notation to define what is called the default alternative, denoted by $a^{*}$. The default can be interpreted as abstention or failing to make a choice. The idea of default is pretty common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^4] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}$, $A^{*}=A\cup\{a^{*}\}$.
@@
Here $\mu(D,A)$ can be interpreted as the probability that $D$ subset of $A$ is the consideration set when the menu $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ is supposed to be a consideration set with some positive probability (full support assumption). This definition is looser than Cattaneo et al. (2020) since these authors, in contrast to the model presented below, impose the condition of monotonicity on the attention function $\mu$. A random attention model (RAM) is then defined as follows.
@@
In other words, a random attention model is a stochastic choice function where for each menu, possible subsets of a menu (consideration sets) are considered/paid attention to with some positive probability. The probability that an alternative is chosen from a menu is the sum of the attention probabilities of consideration sets where that alternative is the most preferred with respect to the underlying preferences.
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that RAM without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set.
Now I will introduce the Random Attention Model with Unobserved References. Here I call the references to be exactly those alternatives that are attention privileged.
I start by first introducing a reference-dependent attention function, it being defined with respect to a pre-specified set $E\subseteq X$ of reference alternatives.
@@
A reference-dependent attention function with the set of references being $E$ is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
@@
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^6] This is the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability..
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly the definition of a reference alternative in this setting. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by these authors since their primitives are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$. This allows us to interpret the default as a reference as well.
The fundamental analytical task in my model is to conduct a revealed preference exercise, i.e. identify the set of reference alternatives and the underlying preferences. In most canonical papers, the primitives assume observability of the references, which precludes the need to identify the set of references.
@@
2.2 Characterization
I characterize the RAM-UR by utilizing three testable axioms. When we are dealing with stochastic choice functions axioms usually restrict how the choices behave across different menus.
@@ 2.2 Characterization
The idea captured by EDA is that of full support. The random process that determines the attention paid to various alternatives allows all alternatives to be considered with positive probability, so the only way an alternative can be never chosen from a menu is if whenever it is considered, it is dominated by another alternative which is considered. An alternative that is attention-privileged and preferred to it will certainly dominate it. It is also a necessary condition because of two properties that the full support assumption implies. First, it is not possible that an alternative is never chosen because it is never considered since all possible supersets of $A\cap E$ have a positive probability. Secondly, if there is no dominant alternative in the menu, there will be a subset of the menu with positive attention probability where the alternative will be the best.
@@
Certainty WARP captures the notion that if an alternative is revealed to be dominated by an attention-privileged alternative, then it will never be chosen with a positive probability whenever that alternative is present. I call it C-WARP because it captures such revealed preference only when an alternative is chosen with certainty from a menu.
@@
Expansion captures the idea that as long as we keep observing certain choices in menus, there exists an attention-privileged alternative that dominates each alternative in the add of these menus. If we look at two attention-privileged alternatives, they will both be always considered, so the preference will determine which will be chosen, but it will be chosen with certainty in the union. It is in some ways similar to the expansion condition of Manzini and Mariotti (2007) in that it deals with the union of two menus. However, it works with any two menus with certain choices. It is silent on menus where certain choice is not observed.
My first result stated below shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
@@
Even with limited data about choices from menus, in most cases we can identify the set of references uniquely. For instance, if there is no data available on singleton sets but we have the choice probabilities from all two and three-element menus, identification will be mostly possible. A rare exception is the case where the second least preferred alternative is not a reference but the least preferred alternative is. In such a case, we will not be able to ascertain which of the two will be the least preferred (and hence, a reference) and which is the second least preferred. However, for other alternatives, we will be able to identify uniquely.
@@
The most important identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives and the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^7] A strict partial order is transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
@@
Kovach and Suleymanov (2023) were able to identify the complete underlying preferences from the general model with observed references. The RAM-UR works with stochastic choice functions and the references are unobservable. Given the significant restrictions to the primitives in this model it is natural that the preference identification result would be weaker. However, even in the restricted model, quite a significant part of the preferences are identified. Even among non-reference alternatives, such that they don’t have the same relative ranking with respect to the references, unique identification of the preferences among them is achieved.
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. In doing so, I characterize the RAM-UR-IRA by restricting the attention function to be of the Independent Random Attention (IRA) form as in Manzini and Mariotti (2014).

### v3

@@ Abstract
In this paper, I develop and characterize a random attention model with unobserved reference alternatives. The decision-maker pays attention to different subsets of the available set of alternatives randomly. The reference alternatives are exactly those alternatives that are always paid attention to, i.e. they are attention-privileged. These alternatives are unknown to the outside observer. The characterization allows for a complete identification of the reference alternatives and a coarse identification of the underlying preferences. I then restrict the model by considering the independent random attention function and provide a complete identification of the underlying preferences.
@@ 1 Introduction
Decision-making is often constrained by limited attention, whether due to cognitive processing limits caused by the availability of too many alternatives or a shortage of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times with the same underlying primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices, with the main research task then being to ascertain the likely underlying heuristic and the primitives that drive the decisions of the DM. Cattaneo et al. (2020) introduced the Random Attention Model (RAM) framework to deal with the analysis of stochastic choice functions with random choice due to limited attention.
@@
A substantial body of literature has explored how an observable reference affects the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. They were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. It is reasonable to assume that a reference alternative can direct attention when the underlying heuristic is applied due to limited attention. They provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
@@
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, their current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus and thus, we do not have to deal with families of stochastic choice functions for each possible reference.
@@
It is also not obvious that a reference alternative must be unique. In what follows, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. Since I am looking at choices from menus that depend on unobservable reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below three examples that match the mentioned assumptions.
  • Example 1: Michael Jordan is associated with the Air Jordan brand. A fan of his who is interested in buying shoes might consider an Air Jordan shoe as a reference. An outside observer need not know that the Air Jordan is in fact a reference for the DM. It is also likely that it stays as a reference whenever the DM has to make a shoe choice, so it is menu-independent. Finally, it does not preclude the DM to also be a fan of another athlete which allows this DM to have multiple reference alternatives.
@@
  • Example 2: Suppose your neighbor buys a new flat-screen 4K TV and invites you over to watch a football game together. You are impressed by the picture quality and ask about the product. The prior knowledge of this particular TV will mean that this product will act as a reference point the next time you buy a TV. Knowledge of such a reference to the outside observer is unlikely, and it also keeps the possibility of multiple references intact.
  • Example 3444I would like to thank Marco Mariotti for this application of the model.: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. The outside observer only observes the current feasible sets and the choices from those but not the choices in the past. In this case, (multiple) past choices act as references and are thus, unobservable to the researcher.
@@
In this paper, I look at a random attention model with unobservable menu-independent reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the RAM-UR by using three axioms and identifying the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences). One can uniquely identify the set of references but the underlying preferences are only coarsely identified. The identification of the preferences gives the relative ranking between any alternative and any reference alternative and the transitive closure thereof. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be fully identified. The existence of menus with no references is crucial in making the preference identification in this case. This is possible because of the menu-independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive which precludes them from considering menus without references. I also elaborate on the fact that rational choice is a boundary case of both the RAM-UR and the RAM-UR-IRA, while the IRA model is itself a boundary case of the RAM-UR-IRA [^5] Boundary cases can be of two types: When the set of references is the set of all alternatives, or when the set of references is empty..
@@
The rest of the paper is organized as follows. In Section 2, I introduce the general RAM-UR model, the axiomatic characterization of the model is presented in Section 3 along with a discussion on the identification results. The introduction and characterization of the RAM-UR-IRA are then provided in Section 4, where I also discuss the nestedness properties of the corresponding random attention models. Section 5 relates my work to the existing literature. All proofs are relegated to Appendix 1.
@@ 2 Random Attention Model with Unobserved References
Below I denote by $a^{*}$ the default alternative and interpret it as abstention or failing to make a choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^6] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}$, $A^{*}=A\cup\{a^{*}\}$.
@@
Here $\mu(D,A)$ can be interpreted as the probability that $D$ is the consideration set when the menu offered is $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ can be a consideration set with some positive probability (full support assumption). In contrast to the above definition, Cattaneo et al. (2020) impose a monotonicity condition on the attention function $\mu$ and do not impose the full support assumption. Kovach and Suleymanov (2023) imposed the full support assumption and it will be crucial in our model as well to identify the set of references. Without the full support assumption, we cannot differentiate between the reference and some of the non-reference alternatives. I also relax the monotonicity condition in the general model. As demonstrated later, this relaxation does not result in much loss in identification. A random attention model (RAM) is then defined as follows.
@@
In other words, a random attention model is a stochastic choice function where for each menu, each possible subset of a menu (consideration sets) is considered/paid attention to with some positive probability. The probability that an alternative is chosen from a menu is the sum of the attention probabilities of consideration sets where that alternative is the most preferred with respect to the underlying preferences.
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that the RAM without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set. The empirical content of the RAM in Cattaneo et al. (2020) comes from the monotonicity restriction they impose on the attention function.
Below I introduce the Random Attention Model with Unobserved References with references being exactly those alternatives that are attention privileged.
I start by first introducing the notion of a reference-dependent attention function.
@@
Let $E\subseteq X$ be a set of reference alternatives. A reference-dependent attention function with respect to $E$ is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
@@
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^8] This is the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability.
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly how I define a reference alternative in this setting. A reference alternative can have other properties, which maybe correlated with the preferences, dependent upon some status quo, or some attribute (the color of a car). However, the presented model is silent on any such dependence. An alternative is a reference if and only if it is attention privileged, irrespective of how it attained such a privilege. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) since the primitives there are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as a reference.
The fundamental analytical task with respect to the presented model is to identify the set of reference alternatives and the underlying preferences. In most canonical papers, it is assumed that the references are observable, which precludes the need to identify the set of references.
@@
3 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict how the choices behave across different menus.
@@ 3 Characterization
A DM with limited attention fails to choose an alternative in two cases. First, if that alternative is never considered. Second, if whenever it is considered, it is dominated by another alternative. EDA says that, the only way an alternative is never chosen is if another alternative always dominates it and is always considered in the binary choice between the two. This means that it is not possible for an alternative to never be chosen just because it is never considered, making each alternative to be non-trivial in the problem. The axiom also says that the set of possible consideration sets is rich enough that the only way an alternative is never chosen is if another dominating alternative is always considered, i.e. it is attention privileged. This way, EDA captures the essence of the modified full support assumption.
@@
Certainty WARP captures the idea that if an alternative is revealed to be dominated by an attention-privileged alternative, then it will never be chosen with a positive probability whenever that alternative is present. I call it C-WARP because it captures such revealed preference only when an alternative is chosen with certainty from a menu.
@@
Expansion captures the idea that as long as we keep observing certain choices in menus, there exists an attention-privileged alternative that dominates each alternative in the union of these menus. If we look at two attention-privileged alternatives, they will both be always considered, so the preference will determine which will be chosen, but it will be chosen with certainty in the union. It is in some ways similar to the expansion condition of Manzini and Mariotti (2007) in that it deals with the union of two menus. However, it works with any two menus with certain choices. It is silent on menus where certain choice is not observed.
The three axioms are independent and restrict the stochastic functions only where deterministic choice or non-choice is observed. This is because the model is very flexible on the structure of the attention function, once the full support assumption is satisfied. To further restrict the stochastic function some regularity condition similar to Cattaneo et al. (2020) must be imposed. However, such an imposition is not necessary to get the first set of results described below.
My first result shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
@@
Even with limited data about choices from menus, in most cases we can identify the set of references uniquely. For instance, if there is no data available on singleton sets but we have the choice probabilities from all two and three-element menus, identification will be mostly possible. A rare exception is the case where the second least preferred alternative is not a reference but the least preferred alternative is. In such a case, we will not be able to ascertain which of the two will be the least preferred (and hence, a reference) and which is the second least preferred. However, we will be able to identify all other reference alternatives uniquely.
@@
The most important identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives and the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus, one can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^9] A strict partial order is transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
@@
Kovach and Suleymanov (2023) were able to identify the complete underlying preferences from the general model with observed references. The RAM-UR works with stochastic choice functions and the references are unobservable. Given the significant restrictions to the primitives in this model, it is natural that the preference identification result is weaker. However, even in the restricted model, quite a significant part of the preferences are identified. Even among non-reference alternatives, with different relative ranking with respect to the references, unique identification of the preferences among them is achieved.
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. First, I will show that the monotonicity condition imposed by Cattaneo et al. (2020) does not offer us any progress in the identification results.
Definition 6
A reference-dependent monotonic attention function with respect to a set $E\subseteq X$ of references is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu_{E}(D,A)=1$
  2. 2.
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$
  3. 3.
    For $S,T\in\mathcal{X}$, if $a\in S\setminus T$, then $\mu_{E}(T,S\setminus\{a\})\geq\mu_{E}(T,S)$
The example below illustrates that even if we restrict ourselves to monotonic attention functions, a unique identification of the preferences is not always possible.
Example 7
Suppose we have three alternatives $X=\{a,b,c\}$ and the following observed choice probabilities:
$p(b,\{a,b\})=1/2\hskip 14.22636ptp(c,\{a,c\})=1/2\hskip 14.22636ptp(a,\{a,b\})=p(a,\{a,c\})=1/2$
$p(a,\{a,b,c\})=p(c\{a,b,c\})=1/4\hskip 14.22636ptp(b,\{a,b,c\})=1/2$
$p(a,\{a\})=1\hskip 14.22636ptp(b,\{b\})=p(c,\{c\})=1/2$
$p(b,\{b,c\})=1/4\hskip 14.22636ptp(c,\{b,c\})=1/2$
We can rationalize these observed choice probabilities through two reference-dependent monotonic attention functions. In both cases $E=\{a\}$ and both $b$ and $c$ are revealed preferred to $a$.
  • $\mu^{1}_{E}$ is the function such that attention is divided among all eligible [^10] Supersets of {a} in any menu containing $a$, and all subsets in any other menu. sets in each menu with the preference being $b\succ c\succ a$.
  • Construct $\mu^{2}_{E}$ as follows:
    $\mu^{2}_{E}(\{a,b\},\{a,b,c\})=1/2\hskip 14.22636pt\mu^{2}_{E}(\{a\},\{a,b,c\})=1/4$
    $\mu^{2}_{E}(\{a,c\},\{a,b,c\})=\mu^{2}_{E}(\{a,b,c\},\{a,b,c\})=1/8$
    $\mu^{2}_{E}(\{a,b\},\{a,b\})=1/2\hskip 14.22636pt\mu^{2}_{E}(\{a,c\},\{a,c\})=1/2$
    $\mu^{2}_{E}(\{b\},\{b,c\})=1/8\hskip 14.22636pt\mu^{2}_{E}(\{b,c\},\{b,c\})=1/8\hskip 14.22636pt\mu^{2}_{E}(\{c\},\{b,c\})=1/2$
[…]
    With the above attention function which is also monotonic and the preference $c\succ b\succ a$ we can also rationalize the observed choice probabilities.
Monotonicity in itself is not enough to refine our identification results, however, further restrictions will be. In this direction, I characterize the RAM-UR-IRA by restricting the attention function to be of the Independent Random Attention (IRA) form as in Manzini and Mariotti (2014). The RAM-UR-IRA then allows for a unique identification of the underlying preferences and the attention function.